# Bibliography management: biblatex package

### v2

@@ 3 Probability
One approach to determining the acceptability of an increase in risk associated with the impact of AI on a given step in the scenario could be to qualitatively evaluate the relative contribution of each step to the overall scenario and then compare this evaluation with the potential benefits of the AI technology. In the biorisk chain, each step has different limiting factors making the relative contribution of $p_{i}$ to $P_{i}$ variable (Table 2). Given that the magnitude of the risk is not known, experts are often relied upon to conceptualize a meaning of “low/med/high”. This could then allow for an assessment of if the impact of AI on $s_{i}$ results in an $s_{ai}$ with an increased or decreased $p_{ai}$ compared to $p_{i}$, potentially changing the categorical contribution to $S_{i}$.
@@
Table 2: Notional requirement levels for completion of each step within the biorisk chain and associated probability. Note: The entries in this table are judgements made by the author alone. It is not intended to serve as an authoritative source on requirement levels for each step.
Step Time Cost Knowledge Resources Safeguard $p_{i}^{step}$
Insight Low Low Low Low Low High
@@
Sequence Low Low Low Low Med High
Biosystem High High High High High Low
Testing High High High High High Low
Dispersal Low High High Med High Med
@@
In this biorisk paradigm, concerning impacts of AI would result in changes of $p_{ai}$ from “low” to “medium” or “high” and from “medium” to “high”. Given the notional content of Table 2, impacts of AI on the biosystem, testing, and dispersal steps may be more important to understand and analyze than those at the insight, ideation, or sequence steps.
4 Consequence
The above scenario comparison is somewhat tractable because the consequence term remains fixed. An increase in the consequence term would also result in an increase in biosecurity risk, assuming that the probability term does not decrease commensurately or more. However, such an assumption is likely unwarranted given the current state of AI technologies in the life sciences and biological engineering. A commonly stated concern is that AI tools may be able to design pathogens that are more transmissible and more pathogenic than those found in nature ([13]). These characteristics of a pathogen are often used as a proxy for consequence. However, as alluded to in the salad bar scenario above, the public health response may differ depending on the severity of disease. Therefore, assessments about the consequence of a deliberate biological attack should be accompanied by a significant amount of uncertainty because it is dependent on unknowable human behavior. Furthermore, this scenario assumes that AI-assisted engineering methods will be able to reach a set of viable end-states (e.g., pathogen design) that are not included within the set of viable end-states of traditional biological engineering methods. To date, this assumption remains largely unproven and lacks adequate rationale. Given the ability to explicitly synthesize DNA sequences, there is no reason to think that an AI tool would generate a design that could not theoretically be designed or identified with sufficient rounds of traditional engineering, such as random mutagenesis. For comparison purposes, and until proven otherwise, it is therefore likely better to assume that traditional engineering approaches and AI-assisted engineering approaches are able to reach an outcome of similar consequence.
If, for the sake of comparison, it is assumed that both $S_{i}$ and $S_{ai}$ have the same set of consequences, the impact of purpose-built AI-based engineering tools on the $p_{ai}$ of each $s_{ai}$ compared to the $p_{i}$ of each $s_{i}$ should be explored, similar to the evaluation of the impact of the chatbots in the prior example. Consider a scenario in which a nefarious actor wants to engineer a pathogen using existing bioengineering methods and the paired scenario containing the use of AI. In this scenario pair, it is assumed that the use of genetic engineering within the biosystem step that requires a traditional synthetic biology process of design – build – test – learn, with the end goal of optimizing the characteristic properties of a pathogen (not the pathogen’s effects on its host) is used ([8]). There is a biomolecule, or “candidate sequence”, that has desirable properties that could be optimized. In $s_{i}^{seq}$ it is assumed the use of random mutagenesis or equivalent method to identify the variants of the candidate sequence to build and test, whereas in $s_{ai}^{seq}$ an AI system generates the recommended sequences to build and test. For the purposes of this evaluation, it is assumed that the requirements to perform mutagenesis and AI design are equal(although, in practice this is likely not true). In $S_{i}$ and $S_{ai}$, $s^{build}$ and $s^{test}$ are identical because they are independent of the AI tool. It is also assumed the $s^{learn}$ is equal between $S_{i}$ and $S_{ai}$, although the specific process of learning from data would proceed differently in each. Therefore, the contribution of AI comes from potentially decreasing the overall number of design-build-test-learn cycles which would reduce the time, cost, and resources required to achieve the desired end goal. To support risk analysis, an understanding of the difference in number of design-build-test-learn cycles between a typical $S_{i}$ and $S_{ai}$ (as a means to characterize a difference in probability of success, with decreasing probability associated with increasing numbers of cycles) is required.
Currently, there is not enough information in the published literature to support a reasonable, generalized characterization of this difference. Often, when researchers are developing AI systems and tools for engineering biology, they will perform the same number of design-build-test-learn cycles to demonstrate that the AI method is improved over the traditional engineering method. For example, a previously reported machine learning method provided a nearly 29-fold improvement compared to a traditional method ([10]). However, it took the team 12+ months to develop the machine learning method that provided the designs. If accounting for the time (and cost, knowledge, resources, etc.) a more apt comparison may be to the number of traditional engineering cycles that could have occurred over that 12+ month development period. Additionally, the authors tested their methods in two very similar engineering tasks and notably, the method that resulted in our best performance in the first scenario gave the worst performance in the second scenario without improving over the traditional engineering method. And lastly, the authors did not perform additional traditional engineering cycles to determine the number of cycles it would take to reach performance of the best AI outcome. To fully understand the impact of including AI methods and tools in biological engineering processes in a manner that supports risk assessment, dedicated efforts to characterize the performance of AI methods across diverse tasks is needed.
Efforts to perform characterization of the ability of AI tools to assist in the engineering pathogens are likely to be controversial. Given that the output of an AI-enabled bioengineering tool is a prediction, validation of the prediction in a biological laboratory is necessary to determine how often the predictions are correct and if there are circumstances or conditions, such as amount or type of training data, that influence the rate of success. Unlike chemical threats, biological threats can transmit from person-to-person and self-amplify. As such, a small mistake in a laboratory could result in large-scale consequence. With this in mind, there should be limits to the extent of characterization of the ability of AI tools to assist in the engineering of pathogens for the sole purpose of informing risk analysis. Creating and evaluating biological agents that are predicted to be highly pathogenic should not be done for the sole purpose of characterizing the predictive capabilities of these tools. This should be the norm, and again, it should not be expected that a quantitative understanding of the impact of AI on biorisk to be achievable.
Proxy evaluation metrics for the ability of AI tools to deliberately engineer pathogens could offer a possible path forward. However, it is unlikely the biosecurity and associated policy communities will be able to agree on sufficient proxy metrics given longstanding disagreements about the ability to “reasonably anticipate” the outcome of scientific experiments ([5]). Therefore, it may be better to consider the capability of an AI tool as a clearly stated assumption in risk analysis conversations or to seek a relative understanding of the level of concern associated with capabilities and scenarios in which biology is used to cause harm. However, the notable number of assumptions required to use the triplet framework described herein may limit its utility in quantitative assessments.
5 A Path Forward
@@
    Identify organisms of interest. Organisms may biologically differ in the way in which their traits of interest are determined. As such, the way these traits are engineered may be different and thus the AI-enabled tools used to support engineering may be different. This list of organisms could include pathogens with pandemic potential, those that are likely candidates for misuse (e.g., Category A), and those that are often engineered in beneficial biotechnological applications because they likely have existing AI-enabled engineering tools.
@@
    Identify existing AI-enabled tools that are amenable to engineering the traits of the organisms of interest through literature and web search. Some tools may be purpose-built for engineering whereas other tools may be designed to characterize organisms and in some scenarios could likely be repurposed or modified to engineer the characterized trait.
@@
Similar processes could also be developed and executed for assessing the impact of AI on the capabilities that do not include the engineering of pathogens, such as those that modify the human microbiome, immune system or genome. In aggregate, the outcome of this potential work would be a relative comparison of levels of concern about the capabilities. It would not necessarily allow for a conclusion about overall changes in the absolute level of biorisk, but rather an understanding of how AI impacts potentially concerning scenarios in a relative manner. Even without an absolute quantitative understanding of biorisk, a relative helps risk managers devote the appropriate attention and resources to mitigate biorisk.
As AI-enabled biological engineering tools continue to mature, it will be increasingly important to understand the ways in which they impact the deliberate biological threat landscape. Such risk analysis efforts should be mindful to ensure that the results are useful to risk managers. Given the complexity of the misuse of biology, risk managers should expect and be supportive of relative, qualitative risk analysis that is informed by a technical understanding of the ways in which AI tools are used within biology and the life sciences. This technical understanding can be furthered by dedicated characterization efforts of AI-enabled tools outside of pathogen engineering context. This work is noticeably absent at present.

### v3

@@ 5 Qualitative Risk
An alternative approach to determining the acceptability of an increase in risk associated with the impact of AI on a given step in the scenario could be to qualitatively evaluate the relative contribution of each step to the overall scenario and then compare this evaluation with the potential benefits of the AI technology. In the biorisk chain, each step has different limiting factors making the relative contribution of $p_{i}$ to $P_{i}$ variable (Table 2). Given that the magnitude of the risk is not known, experts are often relied upon to conceptualize a meaning of “low/med/high”. This could then allow for an assessment of if the impact of AI on $s_{i}$ results in an $s_{ai}$ with an increased or decreased $p_{ai}$ compared to $p_{i}$, potentially changing the categorical contribution to $S_{i}$.
@@
Table 3: Notional requirement levels for completion of each step within the biorisk chain by an unskilled actor and associated probability. Note: The entries in this table are judgements made by the author alone. It is not intended to serve as an authoritative source on requirement levels for each step.
Step Time Cost Knowledge Resources Safeguard Relative $p_{i}^{step}$
@@
Acquisition Low Low Low Low Med Low
Production High High High High High Low
Weaponization High High High High High Low
Deploy/Delivery Low High High Med High Med
@@
In this biorisk paradigm, concerning impacts of AI would likely result in changes of $p_{ai}$ from “low” to “medium” or “high” and from “medium” to “high”. Given the notional content of Table 2, impacts of AI on the acquisition, production, and weaponization steps may be more important to understand and analyze than those at the ideation step.
@@
    Identify organisms of interest. Organisms may biologically differ in the way in which their traits of interest are determined. As such, the way these traits are engineered may be different and thus the AI-enabled tools used to support engineering may be different. This list of organisms could include pathogens with pandemic potential, those that are likely candidates for misuse (e.g., Australia Group List), and those that are often engineered in beneficial biotechnological applications because they likely have existing AI-enabled engineering tools.
@@
    Identify existing AI-enabled tools that are amenable to engineering the traits of the organisms of interest through literature, web search, and key informant interviews. Some tools may be purpose-built for engineering whereas other tools may be designed to characterize organisms and in some scenarios could be repurposed or modified to engineer the characterized trait.
@@
Similar processes could also be developed and executed for assessing the impact of AI on the capabilities that do not include the engineering of pathogens, such as those that modify the human microbiome, immune system or genome. In aggregate, the outcome of this potential work would be a relative comparison of levels of concern about the capabilities. It would not necessarily allow for a conclusion about overall changes in the absolute level of biorisk, but rather an understanding of how AI impacts potentially concerning scenarios in a relative manner. Even without an absolute quantitative understanding of biorisk, a relative understanding helps risk managers devote the appropriate attention and resources to mitigate biorisk.
6 Moving Forward
As AI-enabled biological engineering tools continue to mature, it will be increasingly important to understand the ways in which they impact the deliberate biological threat landscape. Such advances include retrieval-augmented generation (a technique that supplements an LLM’s internal knowledge with external, predefined sources), the development and release of multi-modal models, and the ability of LLMs to access the internet and other software tools, including other AI tools purpose built for life science applications. ([lala_paperqa_2023, bran_chemcrow_2023]). These ongoing efforts have started to blur the distinction between LLM-based chatbots and purpose-built AI tools for engineering biology. Methodologies for evaluating the impacts of AI on all steps of the biorisk chain will need to keep pace with these advances and likely require dedicated laboratory resources.
Given the complexity of the misuse of biology, risk managers and risk assessors will need to work together to clearly articulate their goals and assumptions within their work. Quantitative and qualitative risk analysis, informed by a technical understanding of the ways in which AI tools are used within biology and the life sciences, will be useful, each with different limitations and benefits. The technical understanding of how AI is used and performs within varied life science scenarios can be furthered by dedicated characterization efforts of AI-enabled tools outside of pathogen engineering context. This work, however, is noticeably absent at present.